a chemist has two solutions of the same acid, one containing 15% of the acid and the other 25%. how many cubic centimeters of each should he use to make 120 cc of a 21% acid solution?
Answer by mananth(16946)

a chemist has two solutions of the same acid, one containing 15% of the acid and the other 25%. how many cubic centimeters of each should he use to make 120 cc of a 21% acid solution?
15%.................x cc
25%.................120-x cc
21%................. 120 cc
..
15x+25(120-x)=21*120
15x+3000-25x=2520
-10x=-480
/-10
x=48 cc of 15%
120-x====>120-48=72 cc
..
m.ananth@hotmail.ca